how many liters of water should be added to 5L of a 60% salt solution to produce a solution that is no less than 30% salt and no more than 40% salt?
I set it up as
.30 <= (5)(.60)+(0)(x) <= .40

Answer by Alan3354(69443)

how many liters of water should be added to 5L of a 60% salt solution to produce a solution that is no less than 30% salt and no more than 40% salt?
I set it up as
.30 <= (5)(.60)+(0)(x) <= .40 --> .30 <= .30 <= .40 ??
---------------
For 30%
5*60 + x*0 = (5+x)*30
---
For 40%
5*60 + x*0 = (5+x)*40